Hi
a florist arranges orchids in bunches of 20 and roses in bunches of 10.she sells each orchid bunch for $25 and each rose bunch for $30 .she earnes a total of $390 if she sells 4 times as many orchid bunches as roses bunch es how many flowers did sell altogether.
Thanks

Answer by Theo(13342)

x = number of orchid bunches.
y = number of rose bunches.

each orchid bunch earns 25 dollars and each rose bunch earns 30 dollars and she earns a total of 390, therefore:

25x + 30y = 390

she sells 4 times as many orchid bunches as rose bunches, therefore:

x = 4y

replace x with 4y in 25x + 30y = 390 to get:

25 * (4y) + 30y = 390

simplify to get 100y + 30y = 390

combine like terms to get 130y = 390

solve for y to get y = 390 / 130 = 3

if y = 3, then x = 4y must be equal to 12.

she sold 12 orchid bunches and 3 rose bunches.

12 * 25 = 300
3 * 30 = 90
total is 390.

the question asked how many flowers did she sell.

each orchid bunch contains 20 flowers and each rose bunch contains 10 flowers.

12 * 20 + 3 * 10 = 240 + 30 = 270 flowers total.

i believe that's what they are asking for.